idekanymore14:

I just need help solving it . I keep getting a decimal . (I will post the ss after I post this)

idekanymore14:

1 attachment
AZ:

So it sounds like you got the proofs part down but you just want to know how to solve it? This is what we're given 5x + 14 - 3x = 10x - 28 - 12x To solve it, we need to get `x` on one side, and everything else on the other side

AZ:

So first let's start by simplifying Remember that we can only add like terms So numbers can be added or subtracted with other numbers If anything has an `x` attached to it, we can only add or subtract other things that also have that `x`

idekanymore14:

so 5x-3x and 10x-12x ?

AZ:

Yes! What is 5x - 3x Basically just forget the x is there for a second, and do the normal subtraction. Imagine that x is an apple You have 5 apples and you give away 3 apples. How many apples do you have left?

idekanymore14:

so 2x + 14 = -2x - 28 ?

AZ:

Excellent, yes! Now we want to move anything with `x` in it to one side, and the numbers to another side So 2x + 14 = -2x - 28 what do you get if you add 2x on both sides? (adding 2x on both sides should get rid of it on the right side because it's a -2x)

idekanymore14:

It would be 4x + 14 = -28 I think .

AZ:

Yes!! Very good! Now we have 4x + 14 on the left side We're adding 14. To move it to the other side, we have to do the opposite of adding so what would we do?

idekanymore14:

subtract it to -28 .

AZ:

yes we have to subtract it on both sides so what is -28 - 14 =

idekanymore14:

-42 .

AZ:

good! so now we have 4x = -42 one last thing before we can get the `x` by itself we have 4 times x to get rid of the 4, we have to do the opposite of multiplication

idekanymore14:

-42/4

AZ:

Good! And do the division now

idekanymore14:

but then we get -10.5